OpenStudy (anonymous):

CALC 2 Find the length of the following curve on the given intervals. y=3x-1 from (1,2) to (3,8)

OpenStudy (jziggy):

f(x) = 3x - 1 f'(x) = 3 Arc Length is \[\int\limits_{1}^{3}\sqrt(1 + (3)^2) dx\] \[\sqrt(10)\int\limits_{1}^{3}dx\] \[\sqrt(10)(x) \] evaluated on 1 to 3 \[\sqrt(10)(3-1) = \sqrt(10)(2)\] Final answer:\[2\sqrt(10)\]